S 1-7 Applying the accounting equation LO A1 a. Total assets of Charter Company equal $860,000 and its equity is $500,000. What is the amount of its liabilities? b. Total assets of Martin Marine equal $660,000 and its liabilities and equity amounts are equal to each other. What is the amount of its labilities? What is the amount of its equity? a. 860,000 b. 660,000 500,000 Prev 10, 12 Next 4) 8
Answer #1

a.

Ans:

Formula

Liabilities=total assets-Equity

=860,000-500,000=$360,000

b.

if liabilities and equity amount having equal amounts to each other= i.e in the ratio of 1:1 so total 2

So

Formula

Total Assets/2=$660,000/2=$330,000

Hence

Liabilities =$330,000

Equity=$330,000